You recently accepted an assignment with Oxford Limited as a financial consultant. One of your first
assignments is the analysis of two proposed capital investment projects. Details of initial investments,
after-tax cash flows and average annual profits are represented below:
Year After-tax Cash flows
Initial investments
Cash flows
Average annual profits
1
2
3
4
Project A Project B
(R50 000)
R26 200
R12 200
R12 200
R5 000
R1 400
(R50 000)
R14 200
R14 200
R14 200
R14 200
R1 700
2.1. Which project has a shorter payback period? Motivate your answer by doing the relevant
calculations (answers expressed in years, months, and days).
1
Expert's answer
2020-12-08T02:47:24-0500

Project A

Year Cashflows (Rs) Cumulative cashflows (Rs)

0 (50,000) -

1 26,200 26,200

2 12,200 38,400

3 12,200 50,600

4 5,000 -

Project B

Year Cashflows (Rs) Cumulative cashflows (Rs)

0 (50,000) -

1 14,200 14,200

2 14,200 28,400

3 14,200 42,600

4 14,200 56,800


Payback period for project A

The project gets the R50,000 in the 3rd year. The balance that remains in the 2nd year is R50,000 - R38,400 (the cumulative cashflow). The balance is R11,600 that should come from 3rd year’s cashflow of R12,200.

If R12,200 = 12 months

R11,600 = X

"X = \frac{ R11,600}{ R12,200} \times 12months= 11.41 months"

But, 1 month = 30 days, therefore, 0.41 months = Y days

"Y = \frac{0.41}{1} \times 30 days = 12 days"

Hence, Payback period for project A = 2 years, 11 months, and 12 days.


Payback period for project B

The project gets the R50,000 in the 4th year. The balance that remains in the 3rd year is R50,000 – R42,600 (the cumulative cashflow). The balance is R7,400 that should come from 4thyear’s cashflow of R14,200.

If R14,200 = 12 months

R7,400 = X

"X = \frac{ R7,400}{ R14,200} \times 12 months = 6.25 months"

But, 1 month = 30 days, therefore, 0.25 months = Y days

"Y = \frac{0.25}{1} \times 30 days = 8 days"


Hence, Payback period for project A = 3 years, 6 months, and 8 days.


From the calculations above, project A has the shortest payback period.